OpenStudy (sighn0more):

Can someone please explain to me how to solve a problem like this? Ryan flew from Wiley Post to Ponca City and back. Ryan maintained an average rate of 450 mph going to Ponca City and an average rate of 400 mph returning to Wiley Post. If the actual flying time for the roundtrip was 1 hour, about how far is it from Wiley Post to Ponca City? Round to nearest mile.

OpenStudy (sighn0more):

Thank you!!

OpenStudy (juanpablojr):

no problem :)

OpenStudy (mathstudent55):

Let's call the one way trip distance d. Let's call the time going t. Then the time returning is 1 - t v = d/t d = vt Going: d = 450t Returning: d = 400(1 - t) Since the distances are the same, d = d, then 450t = 400(1 - t) 450t = 400 - 400t 850t = 400 t = 400/850 d = 450t = 450 * 400/850 d = 3600/17 \(\approx\) 211.8 mi